# Distinguishing Neutron Star vs. Low-Mass Black Hole Binaries with Late Inspiral & Postmerger Gravitational Waves — Sensitivity to Transmuted Black Holes and Non-Annihilating Dark Matter

### v2

@@ 2.3 Benchmark Parameters
Cosmic Explorer (CE) 350
Einstein Telescope (ET) 350
Table 3: Benchmark source luminosity distance for each detector configuration analyzed in this work (see Fig. 4).
3 Methods: Distinguishability of LMBH & BNS Mergers
3.1 Fitting Factor & Bayes Factor
The Neyman-Pearson lemma provides that matched filtering is the optimal method to search for GW signals of known shape in the detector data [102]. The detector signal is modeled as a time-series, $s(t)=h(t)+n(t)$, where $h(t)$ is the signal strain and $n(t)$ is the noise in the detector, which is cross-correlated with a template to generate signal to noise ratio (SNR) statistics of a particular event [103, 70]. The optimal SNR (for this case the norm) for a specific frequency range is [103],
$\rho_{\rm opt}=\sqrt{4\int_{f_{\rm min}}^{f_{\rm max}}\frac{|\tilde{h}(f)|^{2}}{S_{n}(f)}df}\equiv\,\rm norm(h)\,$, (3.1)
where $\tilde{h}(f)$ is the Fourier transform (frequency domain) of the time domain data $h(t)$, which is basically the linear combination of two GW polarizations. $S_{n}(f)$ is the one-sided power spectral density (PSD) of the detector [70],
@@ 3.1 Fitting Factor & Bayes Factor
In reality, there remains some mismatch between the template $h_{T}$ and the detector data, which can be further analyzed to infer the parameter values and errors associated with each of them. If the noise is stationary and Gaussian the likelihood is [104],
@@
Our aim is to see whether, for a fixed signal, an LMBH template is favored over a BNS template or vice versa. This comparison can be performed by calculating the ‘Fitting Factor’ FF between two waveforms $h_{1}$ (signal) and $h_{2}$ (template),
${\rm FF}=\max_{\text{templates}}\mathcal{M}$
$=\max_{\text{templates}}2\int_{f_{\rm min}}^{f_{\rm max}}df\frac{\left(h_{1}^{*}(f)h_{2}(f)+h_{1}(f)h_{2}^{*}(f)\right)}{S_{n}(f)\times\sqrt{\langle h_{1}(f)|h_{1}(f)\rangle}\times\sqrt{\langle h_{2}(f)|h_{2}(f)\rangle}}\,.$ (3.5)
This is the match, $\mathcal{M}=\langle\hat{h}_{1}(f)|\hat{h}_{2}(f)\rangle$, maximized over intrinsic parameters of the models [105, 106], and typically used to estimate the maximum overlap achievable between an observed signal (say $h_{1}$) and the best-matching waveform (say $h_{2}$) within a given template bank. It is normalized, which makes it independent of the extrinsic parameters which affect the norms (for our case, mainly $D_{L}$). The $\rm FF$ of exactly same waveforms is unity, and any deviation of the template from the signal will lead to smaller $\rm FF$ values, resulting in the SNR loss in the detector,
$\text{SNR}_{\rm observed}=\rm FF\times\text{SNR}_{\rm opt}\,.$ (3.6)
A reduction in $\rm FF$ greater than 0.03 corresponds $\gtrsim$ 10% loss in the recovered signal amplitude, and hence leads to a significant decrease in the evidential support for the waveform model [107, 108]. The FF is related to the likelihood ratio and can be used for model selection.
The Bayes Factor BF is simply the ratio between two likelihoods, which determines whether hypothesis 1 ($\mathcal{H}_{1}$) is favored over hypothesis 2 ($\mathcal{H}_{2}$). In our case, the functional form of BF can be approximated (using a saddle-point argument) in terms of FF between two models and the $\rho_{\rm opt}$ (Eq. 3.1) [106, 100], as
@@
Here $\mathcal{B}^{1}_{2}$ is the evidence of $\mathcal{H}_{1}$ over $\mathcal{H}_{2}$, given the signal $s$.
We interpret the BF, computed using the FF between BNS merger (say, $\mathcal{H}_{1}$) and LMBH merger waveforms ($\mathcal{H}_{2}$), for a fixed SNR, as the relative likelihood of each model, given a fixed signal. We take the BNS merger waveform (Fig. 2) as a true signal and compare the fit with LMBH merger waveforms (Fig. 3), for each different EoSs (listed in Table 2), and the other specifications are described in Sec. 2.3. As seen from Figs. 1 &2, at lower frequencies (mostly from the inspiral) these two models have a significant resemblance whereas in the high-frequency regime (dominated by the postmerger) they significantly differ with each other. Depending on the EoS, the tidal effects cause the remnant to either go a prompt collapse (smaller $\Lambda$) or live as a stable remnant for some time (larger $\Lambda$) which results in a second peak (at different frequencies) in the waveform unlike the sinusoidal decay of the strain in the case of LMBH mergers [109, 110, 111]. For our analysis we set the frequency range from 5004000500-4000 Hz, where $500\,{\rm Hz}-f_{\rm merge}$(depending on $\kappa_{2}^{T}$, hence EoS) is the inspiral phase and $f_{\rm merge}-4000$ Hz is the postmerger phase. For existing and proposed detectors we start losing sensitivity after 4000 Hz, resulting in no significant contribution from yet higher frequencies (Fig. 4). We use an analytical fit from Ref. [99] to evaluate $f_{\rm merge}$ using
where $\zeta_{3199.8}=\kappa_{2}^{T}+3199.8\left(1-4\nu\right)$, with $\nu$ being the symmetric mass ratio. Table 2 shows the merger frequency for the EoSs used, along with their $\kappa_{2}^{T}$ values.
Since analytically varying all physical parameters for BNS systems is impractical due to the complex microphysical effects and simulation requirements, we restrict our analysis to eight representative EoSs given in Table 2. We report the values in the next subsection for different choices of detectors and explain the trend in the values. Using the FF and $\rho_{\rm opt}$ for the true signal chosen as BNS, we then calculate the Bayes Factor $\mathcal{B}^{\rm BNS}_{\rm LMBH}$, Eq. 3.7, and comment on the Bayesian evidence in the inspiral and the postmerger part of the signal.
3.2 Projected Sensitivity of Planned Detectors
Figure 4: Amplitude Spectral Density (ASD) of the detector-noise, considered in this study. Grey lines denote the dimensionless strain amplitudes $f\tilde{h}(f)$, for the benchmark BNS merger and LMBH merger, with $D_{L}=1$ Mpc. The different detector-noise ASDs are shown as $\sqrt{fS_{n}(f)}$ to keep it dimensionless. $S_{n}(f)$ denotes the one-sided power spectral density of the detector-noise.
The LVK collaboration, comprising the $2^{\rm nd}$ generation, LIGO, Virgo, and KAGRA detectors, has played a foundational role in GW astronomy by enabling the first direct detections of compact binary coalescences and supporting multi-messenger observations [71]. The Neutron Star Extreme Matter Observatory (NEMO), a $2.5$ generation detector, is designed for optimal sensitivity in the high-frequency regime (above $\sim$2 kHz), making it ideal for capturing postmerger signatures from NS mergers and probing the EoS at supranuclear densities [72]. The Cosmic Explorer (CE), a $3^{\rm rd}$ generation detector, aims to achieve ten times the sensitivity of current detectors, expanding the observational horizon to the edge of the observable universe and significantly improving our understanding of GW sources [112, 73]. Similarly the proposed Einstein Telescope (ET), again a $3^{\rm rd}$ generation European detector, is designed to improve sensitivity by an order of magnitude, thereby extending GW detection to cosmological distances and enhancing our ability to probe early-universe phenomena [113, 74]. Together, these detectors cover a broad frequency spectrum: ET and CE excel in the low-frequency band crucial for tracking inspiral dynamics, while NEMO targets the high-frequency postmerger regime, making this combination particularly powerful for distinguishing between NS-NS and LMBH-LMBH mergers and understanding matter effects in extreme gravity.
In Fig. 4 we show the noise sensitivity curve of each of these detectors in the frequency range of our concern. We also show the strain curves for a symmetric BNS and LMBH mergers in dimensionless units. In Table 2 we provide the merger frequency for each different EoS, and for this analysis we take $500\,{\rm Hz}\,-f_{\rm merge}$ as inspiral phase and $f_{\rm merge}-4000$ Hz as postmerger phase. In the following subsections we report our results for each of these detectors and explain how the sensitivity affects the fitting and the Bayes Factor, inferring new details.
3.2.1 Advanced LIGO
EoS norm (Ins) norm (PM) $\rm FF_{Ins}$ $\rm FF_{PM}$ $\rm FF_{total}$ $\rm\mathcal{B}_{Ins}$ $\rm\mathcal{B}_{PM}$ $\rm\mathcal{B}_{total}$
2H 3.51 1.04 0.97 0.77 0.96 1.38 1.24 1.73
MS1b 3.72 0.82 0.98 0.84 0.98 1.26 1.11 1.4
H4 3.89 0.97 0.99 0.71 0.97 1.23 1.26 1.55
@@ 3.2.1 Advanced LIGO
ALF2 4.00 0.75 0.99 0.68 0.98 1.14 1.16 1.33
LS220 2.35 0.76 0.79 0.73 0.77 2.89 1.14 3.42
SLy 4.23 0.62 0.99 0.76 0.99 1.11 1.08 1.21
2B 4.37 0.31 0.997 0.99 0.997 1.06 1.00 1.07
Table 4: Fitting Factor $F$ & Bayes Factor $\mathcal{B}$ for LMBH vs BNS models for different EoS. $\mathcal{B}$ is calculated assuming the BNS as the true signal at LIGO A+ [71]. The $\rho_{\rm opt}$ of each EoS are shown separately for the late inspiral and postmerger phases. The $\mathcal{B}$ for the postmerger part is not enough to draw a conclusion, as the norm in the postmerger part is less than the inspiral part. Here $D_{L}$  is chosen to be 100 Mpc ($m_{1}=m_{2}=1.35\,M_{\odot}$). The EoSs are listed from stiffer to softer cases. Evidently the mismatch is significant in the case of stiffer EoSs. 2B being the softest of all, mimics the LMBH signal, leaving almost equal evidence for both the cases. Interestingly, for the two finite-temperature EoSs - $\rm BHB\Lambda\phi$ & LS220, the fit in the inspiral part is itself worse, resulting in a comparatively higher Bayesian evidence.
Among the mentioned detectors, Advanced LIGO (LIGO A+) has the lowest sensitivity [71], particularly at high and low frequencies, limiting its ability to probe deep inspiral or post-merger dynamics compared to next-generation observatories. In Table 4 we show the norm (or the SNR) of the true LMBH-LMBH signal in both the inspiral and post merger phase, and the Bayes Factor $\left(\mathcal{B}^{\rm LMBH}_{\rm BNS}\right)$ for both the inspiral and postmerger phases, highlighting that for certain stiffer EoSs the Bayesian evidence favoring the LMBH hypothesis is higher in the postmerger phase than in the inspiral phase. Here we chose the $D_{L}$ as 350 Mpc. The BF calculated for the whole frequency range $\mathcal{B}_{\rm total}$ are also shown. BF close to 1 indicates no significant preference between the models. For simplicity of notation, we denote the Bayes Factor $\mathcal{B}^{\rm LMBH}_{\rm BNS}$ as $\mathcal{B}$.
3.2.2 NEMO
EoS norm (Ins) norm (PM) $\rm FF_{Ins}$ $\rm FF_{PM}$ $\rm FF_{total}$ $\rm\mathcal{B}_{Ins}$ $\rm\mathcal{B}_{PM}$ $\rm\mathcal{B}_{total}$
2H 3.06 2.72 0.97 0.71 0.87 1.33 6.32 8.5
MS1b 3.42 2.45 0.98 0.81 0.93 1.27 2.78 3.53
H4 3.67 3.22 0.98 0.7 0.87 1.3 14.8 20.2
@@ 3.2.2 NEMO
ALF2 3.94 2.32 0.99 0.68 0.92 1.22 4.23 5.2
LS220 3.05 2.26 0.9 0.74 0.84 2.4 3.15 8.5
SLy 4.32 1.84 0.99 0.81 0.97 1.2 1.8 2.2
2B 4.71 0.98 0.997 0.99 0.997 1.07 1.01 1.11
Table 5: Fitting Factor $F$ & Bayes Factor $\mathcal{B}$ for LMBH vs BNS models for different EoS with NEMO, taking the BNS merger as the true signal. $D_{L}$ is chosen to be 300 Mpc.
NEMO is expected to operate at cryogenic temperatures and employs shorter, stiffer suspensions to reduce thermal vibrations, especially suspension thermal noise, a dominant source at high frequencies [72]. Other key improvements include the reduction of internal thermal noise from the mirror substrates and coatings, as well as minimizing shot noise, which arises from quantum fluctuations in the laser light [72]. These combined features make NEMO particularly well-suited for capturing the detailed structure of post-merger signals, enabling deeper insight into the NS EoS and other extreme matter phenomena.
In Table 5, we present the values of the FF and $\mathcal{B}$ calculated separately for the inspiral and postmerger phases, as well as for the full frequency range spanning 500 Hz to 4000 Hz. A notable observation is that, owing to NEMO’s enhanced sensitivity in the high-frequency regime, certain stiffer EoSs (H4, 2H, ALF2) yield remarkably higher Bayesian evidence in the postmerger phase compared to the inspiral phase, highlighting the importance of the second peak observed in frequency domain waveform of  Fig.2. As the sensitivity of NEMO is better than LIGO A+, throughout the frequency range of our concern (Fig. 4), we take more distant sources, such as $D_{L}$ = 300 Mpc for an acceptable SNR.
3.2.3 Cosmic Explorer
EoS norm (Ins) norm (PM) $\rm FF_{Ins}$ $\rm FF_{PM}$ $\rm FF_{total}$ $\rm\mathcal{B}_{Ins}$ $\rm\mathcal{B}_{PM}$ $\rm\mathcal{B}_{total}$
2H 14.51 4.11 0.98 0.79 0.97 260.15 24.8 7006.4
MS1b 15.4 3.22 0.99 0.86 0.98 49.5 4.1 210
H4 16.1 3.55 0.99 0.72 0.98 36.7 20.3 747.2
@@ 3.2.3 Cosmic Explorer
ALF2 16.56 2.56 0.99 0.71 0.987 9.8 5.1 53.57
LS220 9.65 2.52 0.79 0.76 0.78 $\mathcal{O}(10^{8})$ 3.75 $\mathcal{O}(10^{9})$
SLy 17.45 2.04 0.994 0.79 0.992 6.01 2.2 13.13
2B 18.04 1.05 0.997 0.988 0.997 2.69 1.01 2.88
Table 6: Fitting Factor $F$ & Bayes Factor $\mathcal{B}$ for LMBH vs BNS models for different EoS with CE, taking the BNS merger as the true signal. The $D_{L}$ is chosen to be 350 Mpc.
Cosmic Explorer (CE) is a proposed third-generation ground-based GW detector designed with 40 km arm lengths, an order of magnitude larger than Advanced LIGO’s 4 km arms, offering vastly improved strain sensitivity across a broad frequency range, particularly in the low-frequency band (510005-1000 Hz), which is crucial for observing the full inspiral phase of CBC with high SNR [73, 112]. In Table 6 we show the FF and $\mathcal{B}$ for each EoS. See the caption of Table 4 for the specifications of the mergers. As the sensitivity of CE significantly exceeds that of LIGO A+, we chose the $D_{L}$ as 350 Mpc. The key observation here is that the Bayesian evidence is already significantly high in the inspiral phase due to CE’s superior low-frequency sensitivity; the inclusion of the postmerger phase further enhances the evidence, offering a more confident characterization of the source.
3.2.4 Einstein Telescope
EoS norm (Ins) norm (PM) $\rm FF_{Ins}$ $\rm FF_{PM}$ $\rm FF_{total}$ $\rm\mathcal{B}_{Ins}$ $\rm\mathcal{B}_{PM}$ $\rm\mathcal{B}_{total}$
2H 14.74 3.67 0.98 0.78 0.97 213.9 14.5 3525.5
MS1b 15.56 2.92 0.99 0.84 0.98 39.19 3.52 147.7
H4 16.22 3.5 0.99 0.72 0.98 28.3 21.4 604.8
@@ 3.2.4 Einstein Telescope
ALF2 16.6 2.6 0.99 0.69 0.986 8.4 6.1 53.13
LS220 9.18 2.58 0.78 0.74 0.77 $\mathcal{O}(10^{8})$ 4.6 $\mathcal{O}(10^{9})$
SLy 17.46 2.15 0.995 0.78 0.992 5.6 2.5 14.2
2B 18 1.08 0.997 0.991 0.997 2.74 1.01 2.93
Table 7: Fitting Factor $F$ & Bayes Factor $\mathcal{B}$ for LMBH vs BNS models for different EoS with ET, taking the BNS merger as the true signal. The source $D_{L}$ is chosen to be 350 Mpc.
The Einstein Telescope (ET) is a proposed third-generation underground GW observatory featuring a unique triangular configuration with 10 km arms, designed to achieve unprecedented sensitivity across a broad frequency range, particularly below 10 Hz, surpassing Advanced LIGO+ and NEMO by mitigating low-frequency seismic, suspension and thermal noise through underground installation and cryogenic technologies [74, 113]. ET, and CE are particularly powerful for studying early inspiral physics, precision cosmology, and multi-band GW astronomy in synergy with space-based detectors. In Table 7 we provide the FF and $\mathcal{B}$ for all the EoSs in both the inspiral and postmerger phase. Here we chose the $D_{L}$ as 350 Mpc. As for CE, for ET the Bayesian evidence is already strong in the inspiral phase. This is primarily due to its excellent low-frequency sensitivity, achieved through underground placement and advanced noise mitigation strategies. The addition of the postmerger phase further improves the evidence.
3.3 Dependence on Luminosity Distance
Having analyzed the FF and the $\mathcal{B}$ across various EoS and for the different detectors, we now investigate how $\mathcal{B}$ varies as a function of the luminosity distance $D_{L}$. We have argued that the waveform fit in the postmerger phase is generally poorer than in the inspiral phase. However, the corresponding Bayesian evidence varies across detectors due to their differing sensitivity profiles. For most cases, the evidence is higher in the inspiral phase, primarily because current detectors exhibit superior sensitivity at lower frequencies. Nonetheless, for stiffer EoS and in detectors with enhanced high-frequency sensitivity; such as NEMO, the Bayesian evidence in the postmerger phase can surpass that of the inspiral phase.

### v3

@@ 2.1 Inputs & Benchmark Parameters
CE & ET 350
Table 1: Benchmark intrinsic parameters, component masses ($m_{1},\,m_{2}$) & dimensionless spins ($\chi_{1},\,\chi_{2}$), used in this study for BNS and BLMBH systems. All sources are assumed to be equal-mass, non-spinning binaries. The benchmark source luminosity distance $D_{L}$ for each detector analyzed in this work is listed alongside.
BNS Microphysics
[…]
Table 2: The EoSs, considered for the BNS systems in this study, with their CoRe IDs. The reduced tidal parameters $\kappa_{2}^{T}$ & $\tilde{\Lambda}$ are also listed (see Eq. A.5 for definitions). The last column gives the predicted merger frequencies [69, 70, 71]. We take the frequency range from 500 Hz to $f_{\rm merge}$ as the inspiral and from $f_{\rm merge}$ to 4000 Hz as the postmerger.
Table 1 lists the intrinsic parameters used for both BNS an BLMBH waveforms. It also notes the benchmark distances $D_{L}$ that we have assumed for sources to be detected at each of the GW detector considered. These benchmark distances are loosely chosen to ensure a sufficiently significant detection in each case, and we show the dependence of our results on $D_{L}$ in Sec. 2.3.2. We also list the EoSs used in this study in Table 2 from stiffer to softer models [69, 72].
2.2 Fitting Factor
The detector signal is modeled as a time-series, $s(t)=h(t)+n(t)$, where $h(t)$ is the signal strain and $n(t)$ is the noise in the detector, which is match-filtered with a bank of templates $h_{T}(t)$ to generate the signal to noise ratio (SNR) statistics of a particular event [73, 74, 58]. The optimal SNR is obtained when $h_{T}(t)=h(t)$, and given as
${\rm SNR_{\rm opt}}\equiv\rho_{\rm opt}=\sqrt{4\int_{f_{\rm min}}^{f_{\rm max}}\frac{|\tilde{h}(f)|^{2}}{S_{n}(f)}df}\equiv\,{\rm norm}(h)\,$, (2.1)
where $\tilde{h}(f)$ is the Fourier transform (frequency domain) of the time domain data $h(t)$ and $S_{n}(f)$ is the one-sided power spectral density (PSD) of the detector [58],
@@ 2.2 Fitting Factor
In reality, there remains some mismatch between the template $h_{T}$ and the detector data, which can be further analyzed to infer the parameter values and errors associated with each of them. If the noise is stationary and Gaussian, the likelihood is [75]
@@
Our aim is to quantify how well a BLMBH merger waveform matches with a given BNS merger waveform. This comparison can be performed by calculating the fitting factor FF between two waveforms $h_{1}$ (signal = BNS) and $h_{2}$ (template = BLMBH),
${\rm FF}=\max_{\text{templates}}\mathcal{M}=\max_{\text{templates}}2\int_{f_{\rm min}}^{f_{\rm max}}df\frac{\left(\tilde{h}_{1}^{*}(f)\tilde{h}_{2}(f)+\tilde{h}_{1}(f)\tilde{h}_{2}^{*}(f)\right)}{S_{n}(f)\times\sqrt{\langle{h}_{1}|{h}_{1}\rangle}\times\sqrt{\langle{h}_{2}|{h}_{2}\rangle}}\,.$
This is the match, $\mathcal{M}=\langle\hat{h}_{1}(f)|\hat{h}_{2}(f)\rangle$, maximized over intrinsic parameters of the models [76, 77], and used to estimate the maximum overlap achievable between an observed signal (say $h_{1}$) and the best-matching waveform within a given template bank (say $h_{2}$). It is normalized, which makes it independent of the extrinsic parameters which affect the norms (for our case, mainly $D_{L}$). The $\rm FF$ of exactly the same waveforms is unity, and any deviation of the template from the signal will lead to smaller $\rm FF$ values, resulting in the SNR loss in the detector,
$\text{SNR}_{\rm observed}=\rm FF\times\text{SNR}_{\rm opt}\,.$ (2.5)
A reduction in $\rm FF$ by more than 0.03 corresponds to a $\gtrsim$ 10% loss in the recovered signal amplitude, and hence leads to a significant decrease in the evidential support for the concerned waveform [78, 79].
For our analysis we consider the frequency range from 5004000500-4000 Hz, where we call the range $500\,{\rm Hz}-f_{\rm merge}$(depending on $\kappa_{2}^{T}$, hence EoS) as the late inspiral phase and $f_{\rm merge}-4000$ Hz as the postmerger phase. The sensitivity drops after 4000 Hz, resulting in no significant contribution from higher frequencies (Fig. 1). We use an analytical fit from Ref. [69] to evaluate $f_{\rm merge}$ using
where $\zeta_{3199.8}=\kappa_{2}^{T}+3199.8\left(1-4\nu\right)$, with $\nu$ being the symmetric mass ratio (see Appendix  A.2 for further details).
Inspiral      Postmerger       Total
Refer to caption Refer to caption Refer to caption
@@
Since analytically varying all physical parameters for BNS systems is impractical due to the simulation costs, we restrict our analysis to eight representative EoSs given in Table 2. Table 2 shows the $\kappa_{2}^{T}$ values and predicted merger frequencies.
In Fig. 2 we show the matches and fitting factors for our chosen suite of EoSs and detectors. One can see that the fitting factor is generically poor (<0.97<0.97, corresponding to a loss $10\%$ of the SNR) in the postmerger signal. However, the norm (or equivalently SNR) is usually larger in the inspiral phase (except for NEMO), and the fitting factor for the total signal more closely matches what would have been obtained from the inspiral signal alone. This illustrates the value of instruments that are capable of detecting the high-frequency postmerger signal to distinguish BNSs and BLMBHs. The key point is that the matter effect, encoded via the EoS, determines a characteristic secondary peak in the postmerger GW waveform, that in principle provides a clear and robust distinction between the two cases.[80, 81, 82]. As we shall see shortly, this is unfortunately difficult to leverage because of the relatively lower sensitivity at higher frequencies for all detectors, except at NEMO.
2.3 Bayes Factor
The Bayes factor (BF) is the ratio of two marginal likelihoods, which determines whether hypothesis 1 ($\mathcal{H}_{1}$ = BNS) is favored over hypothesis 2 ($\mathcal{H}_{2}$=BLMBH). More accurately, $\mathcal{B}^{1}_{2}$ is the evidence of $\mathcal{H}_{1}$ over $\mathcal{H}_{2}$, given the signal $s=h_{\rm BNS}$. In our case, the functional form of the BF can be approximated (using a saddle-point argument) in terms of the FF between the two models (with equal a priori odds) and the optimal SNR $\rho_{\rm opt}$ (Eq. 2.1) [77, 70], as
@@ 2.3 Bayes Factor
We interpret $\mathcal{B}^{\rm BNS}_{\rm BLMBH}\equiv\mathcal{B}$ as the relative likelihood of BNS vs. BLMBH models. We will use Eq. 2.7 to estimate $\mathcal{B}$ — in the late inspiral signal, in the postmerger signal, and in the total signal — at each detector, and for our suite of EoS.
2.3.1 Distinguishability at Current & Proposed Detectors
Advanced LIGO
Distinguishability at LIGO A+
@@ 2.3.1 Distinguishability at Current & Proposed Detectors
Table 3: Fitting factor FF & Bayes factor $\mathcal{B}$ for BNS vs BLMBH merger models, for different EoSs, computed separately in the late inspiral phase, the postmerger phase, and in total. The norm, i.e. $\rho_{\rm opt}$, is shown for the late inspiral and postmerger phases separately; the total norm is obtained by adding these in quadrature. As one can see, despite potential distinguishability due to low FFs in some cases, the BFs are not significant and the BNS and BLMBH models are essentially indistinguishable. $D_{L}$ is chosen to be 100 Mpc.
@@
NEMO
Distinguishability at NEMO
@@
Table 4: Fitting factor FF & Bayes factor $\mathcal{B}$ for BNS vs. BLMBH models at NEMO. For the stiffer EoSs, there is potential for distinguishing BNSs and BLMBHs, with a significant role played by the postmerger signal. $D_{L}$ is chosen to be 300 Mpc.
NEMO, being more sensitive in the postmerger phase than in the inspiral, is expected to capture matter effects more efficiently. Table 4 presents the FFs and BFs. It is promising that at NEMO there is significant evidence to distinguish the models, especially for the stiffer EoSs (H4, 2H, ALF2). Notably, the evidence from the postmerger phase is significant. We have taken a larger benchmark distance, $D_{L}$ = 300 Mpc compared to LIGO A+.
Cosmic Explorer and Einstein Telescope
In Table 5 we show the FF and $\mathcal{B}$ at the Cosmic Explorer (CE) and Einstein Telescope (ET). Here we chose $D_{L}$ as 350 Mpc. As their sensitivities significantly exceed that of LIGO A+ over the entire frequency range, and that of NEMO in the low frequency range, the BFs are much larger. Now the BNS and BLMBH models can be confidently distinguished for all viable EoSs considered (note that the EoS 2B is currently disfavored from NS mass-radius data). A key observation here is that the Bayesian evidence is already significantly high in the inspiral phase due to their superior low-frequency sensitivity; the inclusion of the postmerger phase enhances the evidence but no longer plays as crucial a role as for NEMO. If further parameter degeneracies (spins, eccentricities) are taken into account it is possible though that the postmerger signal may be more important than is indicated here. We expect this because the postmerger signal is less likely to be degenerate with other parameters.
Distinguishability at Cosmic Explorer
@@
Distinguishability at Einstein Telescope
@@
Table 5: Fitting factor FF & Bayes factor $\mathcal{B}$ for BNS vs BLMBH models at CE (upper table) and ET (lower table), with the signal generated using a BNS waveform with the stated EoSs. The large values of $\mathcal{B}$ indicate that BNS and BLMBH models can be confidently distinguished. The source $D_{L}$ is chosen to be 350 Mpc.
2.3.2 Dependence on Luminosity Distance
We now show how the distinguishing ability varies as a function of the distance of the source $D_{L}$. We can readily conclude from Eq. 2.7 that the dependence is given by $\log\mathcal{B}\propto 1/D_{L}^{2}$. It follows from the fact that $\rm FF$ is independent of $D_{L}$ and $\rho_{\rm opt}$ scales as $1/D_{L}$. One can compare the detectors at a common benchmark source distance, if needed, by scaling the BFs as above.
Figure 3: Bayes Factor ($\mathcal{B}$) as a function of $D_{L}$ for different detectors and different EoSs. The dependence is given by $\mathcal{B}\propto\exp(1/D_{L}^{2})$. The evidence against a BLMBH merger over a BNS merger becomes significantly large for stiffer EoSs. Note that the distance range on the horizontal axis varies across panels.